The solution for 4x-9y=10.5 and 7x-2x=6.5
Answer by mananth(16949)

4 x -9 y = 10.5 .............1
7 x -2 y = 6.5 .............2

multiply (1)by 2
Multiply (2) by -9
8 x -18 y = 21
-63 x 18 y = -58.5
Add the two equations
-55 x = -37.5
/ -55
x = 0.68 OR 2/3
plug value of x in (1)
4 x + -9 y = 10.5
2.73 + -9 y = 10.5
-9 y = 10.5 -2.73
-9 y = 7.77
y = -0.86 OR - 6/7